OpenStudy (anonymous):

write the equation in slope-intercept form: 6(x+3y)=36

OpenStudy (anonymous):

i already did this one

OpenStudy (anonymous):

try to solve them on your own now

OpenStudy (anonymous):

i'll get you started : distribute the 6

OpenStudy (anonymous):

well what was the answer lol sorry

OpenStudy (anonymous):

tell me what you get and i'll check you

OpenStudy (anonymous):

i only have 2 more problems left, please just answer them

OpenStudy (anonymous):

y=-1/2x+2

OpenStudy (anonymous):

thankyou!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! ;)

OpenStudy (anonymous):

oooook so you distribute the 6 and then you get 6x + 18y = 36 then move the 6 over : 18y = 36 - 6x then divide the 18